# On the proximal dynamics between integrable and non-integrable members of a generalized Korteweg-de Vries family of equations

### v1

@@ 3. $\bullet$
Figure 4. Left and middle: Snapshots comparing the profiles of the gKdV solution in the case of the power nonlinearity (1.3) with $k=4$ (solid blue) against the profiles of the soliton (4.7) of the rescaled integrable KdV (4.6) with $\nu=0.41$ (red dashed). Right: The top panel shows the contour plot of the evolution of the initial condition (4.4) with $c=0.4$, $x_{0}=0$ under gKdV. The dashed black line is the caustic of the rescaled soliton (4.7) with $\nu=0.41$. The bottom panel presents the evolution of norms of the distance between the solutions $\mathcal{X}=H_{x}^{s}$, $s=0,1,2$, and $L_{x}^{\infty}$.
@@ 4.2. The weakly perturbed KdV and mKdV equations
I. Single soliton initial data. As a first example, we consider the long-time dynamics of one-soliton solutions. In particular, we simulate numerically the time evolution of the gKdV equation (1.1) with $F(U)=U+\delta U^{3}$ and $\delta=0.02$, supplemented with the KdV one-soliton initial datum $u_{\text{s}}(x)$ in (4.4) with $c=0.3$ and $x_{0}=0$ for which $\left\|u_{\text{s}}\right\|_{H^{2}}\approx 2.05=\epsilon$. While the proximity upper bounds of Theorem 1 exhibit a large growth rate in terms of $\epsilon>1$, a further inspection of the derivation of the relevant estimates (e.g. see inequalities (3.12) and (3.19)) suggests that the overall linear growth rate should be compensated by the smallness of the parameter $\delta$ in the example under consideration. The numerical results, illustrated in Figure 5 over a time interval $t\in[0,T]$ with $T\approx 500$, confirm this compensation. They show strong agreement between the solution dynamics of the integrable KdV equation and its non-integrable perturbation, even at large times, and demonstrate the moderate linear growth of the distances $\left\|\Delta(t)\right\|_{\mathcal{X}}$ for $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$.
Similar numerical results are obtained for the second example, where we simulate the time evolution of the gKdV equation (1.1) with $F(U)=U^{2}+\delta U^{4}$ and $\delta=0.01$. This evolution is initialized via the mKdV one-soliton solution $U_{\text{s}}(x)$ from (4.4) with $c=0.2$ and $x_{0}=0$. For this initial datum, $\left\|U_{\text{s}}\right\|_{H^{2}}\approx 2.41=\epsilon$. The simulation is performed over a time interval $t\in[0,T]$ with $T\approx 500$. As shown in the bottom row of Figure 5, we once again observe strong agreement between the solution dynamics of the integrable mKdV equation and its non-integrable perturbation, even at large times, together with the moderate linear growth of the distances $\left\|\Delta(t)\right\|_{\mathcal{X}}$ for $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$.
@@
where $c_{1}=0.08$, $c_{2}=0.2$, $x_{01}=40$ and $x_{02}=0$. The norm of this initial datum is $\left\|u_{\text{2}s}\right\|_{H^{2}}\approx 0.84=\epsilon$. Since $\epsilon<1$, according to Theorem 1, we expect proximal dynamics to persist for long times. We perform the simulation over a time interval $t\in[0,T]$ with $T\simeq 500$. In agreement with the theoretical analysis, the top rows of Figures 6 and 7 show a remarkable agreement between the soliton collision dynamics of the integrable KdV equation and its non-integrable perturbation. Regarding the evolution of the distances $\left\|\Delta(t)\right\|_{\mathcal{X}}$, it is interesting to note, once again, the appearance of maxima and minima in their time evolution, like in the case of large data for the power nonlinearity (see part II in Section 4.1). These extrema are due to the slight curvature introduced by the weak nonlinearity of the gKdV equation, such that the distance is minimized precisely at the instant of the soliton collision where the caustics of the solitons cross one another.
Figure 6. Top: The solutions of the KdV equation and the gKdV equation (1.1) with $F(U)=U+\delta U^{3}$ and $\delta=0.02$, supplemented with the KdV two-soliton initial datum (4.8) with $c_{1}=0.08$, $c_{2}=0.2$, $x_{01}=40$ and $x_{02}=0$. Bottom: The solution to the mKdV equation and the gKdV equation (1.1) with $F(U)=U^{2}+\delta U^{4}$ and $\delta=0.01$, supplemented with the mKdV two-soliton initial datum (4.9) with $c_{1}=0.2$, $c_{2}=0.1$, $x_{01}=-20$ and $x_{02}=0$.
Figure 7. Top: The solution to the gKdV equation (1.1) with $F(U)=U+\delta U^{3}$ and $\delta=0.02$, supplemented with the KdV two-soliton initial datum (4.8) with $c_{1}=0.08$, $c_{2}=0.2$, $x_{01}=40$ and $x_{02}=0$. Bottom: The solution to the gKdV equation (1.1) with $F(U)=U^{2}+\delta U^{4}$ and $\delta=0.01$, supplemented with the mKdV two-soliton initial datum (4.9) with $c_{1}=0.2$, $c_{2}=0.1$, $x_{01}=-20$ and $x_{02}=0$. Left: The initial profiles. Middle: The solutions at $t=500$. Right: The time evolution of the norms $\left\|\Delta(t)\right\|_{\mathcal{X}}$ with $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$.
@@
The results of the simulation for $t\in[0,T]$ with $T\approx 500$ are depicted in the bottom rows of Figures 6 and 7. Indeed, the third panel of the bottom row of Figure 7, which portrays the evolution of the norms $\left\|\Delta(t)\right\|_{\mathcal{X}}$, shows the divergence of norms at a significantly faster rate than in the case of small initial data of the first example for the weakly perturbed KdV equation. Specifically, comparing with the third panel of the top row of Figure 7, we observe that the growth rate for the large data is of $\mathcal{O}(10^{-1})$. This rate, although notably larger than the rate $\mathcal{O}(10^{-2})$ observed in the case of small initial data, is still moderate due to the small strength of the perturbation $\delta$ of the mKdV equation. This moderate growth still guarantees the close proximity of the collision dynamics between the KdV equation and the perturbed mKdV equation of the example, as shown in the bottom row of Figure 6 and the snapshots of the profiles of the solutions of the bottom row of Figure 7. Returning to the third panel of the bottom row of Figure 7 for the distances $\left\|\Delta(t)\right\|_{\mathcal{X}}$, we again notice the appearance of the extrema, at shorter times. This observation is in accordance with the fact that the weak perturbation $\delta U^{4}$ introduces a larger curvature effect in the caustics of the solitons, similar to the case of $k=4$ for the power nonlinearity (see part II of Section 4.1). The minimum appears again precisely at the instant of the soliton collisions, as explained in the first example for the pertubed KdV equation.
@@ 5. Conclusion
The theoretical and numerical results of the present work pave the way for interesting potential extensions. On the theoretical side, an important direction is the extension of these results to functional space settings appropriate for studying the persistence of spatially periodic solutions, quasi-periodic solutions, or the potential proximal dynamics originating from non-smooth data. Further theoretical directions concern the case of more general nonlinearities, including for example fractional power nonlinearities. Finally, by exploiting complete integrability of the governing model, mathematically precise descriptions of phenomena such as small-dispersion limits, dispersive shock-waves, and soliton gases (random ensembles of interacting solitons) have been obtained. An interesting future direction is the combination of these descriptions with the proximity results of the present work in order to obtain analogous descriptions for non-integrable models.

### v2

@@ 3. $\bullet$
(a)               (b)               (c)
Refer to caption Refer to caption Refer to caption
(d)               (e)               (f)
Refer to caption Refer to caption Refer to caption
Figure 4. (a), (b), (d) and (e) Snapshots comparing the profiles of the gKdV solution in the case of the power nonlinearity (1.3) with $k=4$ (solid blue) against the profiles of the soliton (4.7) of the rescaled integrable KdV (4.6) with $\nu=0.41$ (red dashed). (c) Contour plot of the evolution of the initial condition (4.4) with $c=0.4$, $x_{0}=0$ under gKdV. The dashed black line is the caustic of the rescaled soliton (4.7) with $\nu=0.41$. (f) The evolution of norms of the distance between the solutions $\mathcal{X}=H_{x}^{s}$, $s=0,1,2$, and $L_{x}^{\infty}$.
@@ 4.2. The weakly perturbed KdV and mKdV equations
I. Single soliton initial data. As a first example, we consider the long-time dynamics of one-soliton solutions. In particular, we simulate numerically the time evolution of the gKdV equation (1.1) with $F(U)=U+\delta U^{3}$ and $\delta=0.02$, supplemented with the KdV one-soliton initial datum $u_{\text{s}}(x)$ in (4.4) with $c=0.3$ and $x_{0}=0$ for which $\left\|u_{\text{s}}\right\|_{H^{2}}\approx 2.05=\epsilon$. While the proximity upper bounds of Theorem 1 exhibit a large growth rate in terms of $\epsilon>1$, a further inspection of the derivation of the relevant estimates (e.g. see inequalities (3.12) and (3.19)) suggests that the overall linear growth rate should be compensated by the smallness of the parameter $\delta$ in the example under consideration. The numerical results, illustrated in Figure 5 over a time interval $t\in[0,T]$ with $T\approx 500$, confirm this compensation. Panels (a) and (b) in Figure 5 illustrate snapshots at specific times comparing the profiles of the solutions. They show strong agreement between the solution dynamics of the integrable KdV equation and its non-integrable perturbation, even at large times. Panel (c) demonstrates the moderate linear growth of the distances $\left\|\Delta(t)\right\|_{\mathcal{X}}$ for $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$.
Similar numerical results are obtained for the second example, where we simulate the time evolution of the gKdV equation (1.1) with $F(U)=U^{2}+\delta U^{4}$ and $\delta=0.01$. This evolution is initialized via the mKdV one-soliton solution $U_{\text{s}}(x)$ from (4.4) with $c=0.2$ and $x_{0}=0$. For this initial datum, $\left\|U_{\text{s}}\right\|_{H^{2}}\approx 2.41=\epsilon$. The simulation is performed over a time interval $t\in[0,T]$ with $T\approx 500$. As shown in panels (d) and (e) of Figure 5, we once again observe strong agreement between the solution dynamics of the integrable mKdV equation and its non-integrable perturbation, even at large times, together with the moderate linear growth of the distances $\left\|\Delta(t)\right\|_{\mathcal{X}}$ for $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$, shown in panel (f).
(a)               (b)               (c)
Refer to caption Refer to caption Refer to caption
(d)               (e)               (f)
Refer to caption Refer to caption Refer to caption
@@
where $c_{1}=0.08$, $c_{2}=0.2$, $x_{01}=40$ and $x_{02}=0$. The norm of this initial datum is $\left\|u_{\text{2}s}\right\|_{H^{2}}\approx 0.84=\epsilon$. Since $\epsilon<1$, according to Theorem 1, we expect proximal dynamics to persist for long times. We perform the simulation over a time interval $t\in[0,T]$ with $T\simeq 500$. In Figure 6, panel (a) depicts the contour plot of the analytical two-soliton solution of the KdV equation while panel (b) depicts the contour plot of the solution of the gKdV equation initiated by the above initial datum. For the same example, in Figure 7 panel (a) shows the initial profiles and panel (b) shows the comparison of the profiles at $t=500$. In line with the theoretical analysis, we observe the remarkable agreement between the soliton collision dynamics of the integrable KdV equation and its non-integrable perturbation. Regarding the evolution of the distances $\left\|\Delta(t)\right\|_{\mathcal{X}}$ shown in panel (c) of Figure 7, it is interesting to note, once again, the appearance of maxima and minima in their time evolution, like in the case of large data for the power nonlinearity (see part II in Section 4.1). These extrema are due to the slight curvature introduced by the weak nonlinearity of the gKdV equation, such that the distance is minimized precisely at the instant of the soliton collision where the caustics of the solitons cross one another.
(a)               (b)
Refer to caption Refer to caption
(c)               (d)
Refer to caption Refer to caption
Figure 6. (a) Contour plot of the (analytical) solution of the KdV equation, supplemented with the KdV two-soliton initial datum (4.8) with $c_{1}=0.08$, $c_{2}=0.2$, $x_{01}=40$ and $x_{02}=0$. (b) Contour plot of the solution of the gKdV equation (1.1) with $F(U)=U+\delta U^{3}$ and $\delta=0.02$, supplemented with the KdV two-soliton initial datum (4.8) with $c_{1}=0.08$, $c_{2}=0.2$, $x_{01}=40$ and $x_{02}=0$. (c) Contour plot of the (analytical) solution of the mKdV equation, supplemented with the mKdV two-soliton initial datum (4.9) with $c_{1}=0.2$, $c_{2}=0.1$, $x_{01}=-20$ and $x_{02}=0$. (d) Contour plot of the solution of the gKdV equation (1.1) with $F(U)=U^{2}+\delta U^{4}$ and $\delta=0.01$, supplemented with the mKdV two-soliton initial datum (4.9) with $c_{1}=0.2$, $c_{2}=0.1$, $x_{01}=-20$ and $x_{02}=0$.
(a)               (b)               (c)
Refer to caption Refer to caption Refer to caption
(d)               (e)               (f)
Refer to caption Refer to caption Refer to caption
Figure 7. (a) and (b) Snapshots comparing the profile of the solution of the gKdV equation (1.1) with $F(U)=U+\delta U^{3}$ and $\delta=0.02$, supplemented with the KdV two-soliton initial datum (4.8) with $c_{1}=0.08$, $c_{2}=0.2$, $x_{01}=40$ and $x_{02}=0$, against the profile of the corresponding KdV analytical two-soliton solution. Panel (a) shows the initial profiles and panel (b) the solutions at $t=500$. (c) The corresponding evolution of the norms $\left\|\Delta(t)\right\|_{\mathcal{X}}$ with $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$. (d) and (e) Snapshots comparing the profile of the solution of the gKdV equation (1.1) with $F(U)=U^{2}+\delta U^{4}$ and $\delta=0.01$, supplemented with the mKdV two-soliton initial datum (4.9) with $c_{1}=0.2$, $c_{2}=0.1$, $x_{01}=-20$ and $x_{02}=0$, against the profile of the corresponding mKdV analytical two-soliton solution. Panel (d) shows the initial profiles and panel (e) the solutions at $t=500$. (f) The corresponding time evolution of the norms $\left\|\Delta(t)\right\|_{\mathcal{X}}$ with $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$.
@@
The results of the simulation for $t\in[0,T]$ with $T\approx 500$ are depicted in panels (c) and (d) of Figure 6 and panels (d), (e) and (f) of Figure 7. In Figure 6, panel (c) depicts the contour plot of the corresponding analytical two-soliton solution of the mKdV equation while panel (d) depicts the contour plot of the solution of the gKdV equation initiated by the above initial datum. For the same example, in Figure 7 panel (d) shows the initial profiles and panel (e) shows the comparison of the profiles at $t=500$.
Indeed, panel (f) of Figure 7, which portrays the evolution of the norms $\left\|\Delta(t)\right\|_{\mathcal{X}}$, shows the divergence of norms at a significantly faster rate than in the case of small initial data of the first example for the weakly perturbed KdV equation. Specifically, comparing with panel (c) of Figure 7, we observe that the growth rate for the large data is of $\mathcal{O}(10^{-1})$. This rate, although notably larger than the rate $\mathcal{O}(10^{-2})$ observed in the case of small initial data, is still moderate due to the small strength of the perturbation $\delta$ of the mKdV equation. This moderate growth still guarantees the close proximity of the collision dynamics between the KdV equation and the perturbed mKdV equation of the example, as shown in panels (d) and (e) of Figure 6 and the snapshots of the profiles of the solutions shown in panels (d) and (e) of Figure 7. Returning to panel (f) of Figure 7 for the distances $\left\|\Delta(t)\right\|_{\mathcal{X}}$, we again notice the appearance of the extrema, at shorter times. This observation is in accordance with the fact that the weak perturbation $\delta U^{4}$ introduces a larger curvature effect in the caustics of the solitons, similar to the case of $k=4$ for the power nonlinearity (see part II of Section 4.1). The minimum appears again precisely at the instant of the soliton collisions, as explained in the first example for the pertubed KdV equation.
@@ 5. Conclusion
The theoretical and numerical results of the present work pave the way for interesting potential extensions. On the theoretical side, an important direction is the extension of these results to functional space settings appropriate for studying the persistence of spatially periodic solutions, quasi-periodic solutions, or the potential proximal dynamics originating from non-smooth data. Further theoretical directions concern the case of more general nonlinearities, including for example fractional power nonlinearities. Finally, by exploiting complete integrability of the governing model, mathematically precise descriptions of phenomena such as small-dispersion limits, dispersive shock-waves, and soliton gases (statistical ensembles of interacting solitons) have been obtained. An interesting future direction is the combination of these descriptions with the proximity results of the present work in order to obtain analogous descriptions for non-integrable models.